OpenStudy (anonymous):

x^2+y^2-6x+10y+34=0 is this a circle, a point or no real graph? if A CIRCLE find radius and center

OpenStudy (anonymous):

\[x^2+y^2-6x+10y+34 = (x-3)^2 +(y+5)^2=0 \] It is the point (3,-5)

OpenStudy (anonymous):

ok thanks